# Networks of neural networks: more is different

### v1

@@ 3.2 Stability analysis in the low-load, noisy, and zero-field regime
Figure 3: The solid lines represent the numerical solution of the self-consistency equations (11) in the low-load regime and in the absence of external field, obtained by applying the fixed-point iteration method with initial point given by $\bar{\bm{m}}_{\{\mu\leq L\}}^{(1,2,3)}$ (left) and by $\bar{\bm{m}}_{\{\mu\leq L\}}^{(h)}$ (right), as given in eqs. (25)-(26). These numerical solutions preserve the structure of the initial datum, specifically, on the left, the solid lines show the behavior of $\bar{m}_{1}^{1}=\bar{m}_{2}^{2}=\bar{m}_{3}^{3}$ while $\bar{m}_{\mu}^{a}$ is vanishing for $\mu\neq a$; on the right, the the solid lines show the behavior of $\bar{m}_{\mu}^{a}$, that coincides for any $a\in[1,2,3]$ and $\mu\in[1,2,3]$. The persistency in the structure of the solution is lost at a certain value of $\beta^{-1}$, highlighted by the vertical dotted lines: beyond these values, that depend on $\lambda$ (see the common legend on the right), solutions with a different structure appear, and these correspond, for instance, to the state $\bm{\sigma}^{(1,1,1^{\prime})}$.
Let us now focus on the stability of these solutions: as we will show, $\bm{\sigma}^{(1,2,3)}$ and $\bm{\sigma}^{(h)}$ display different stability curves in the $(\beta,\lambda)$ plane and, in particular, there exists a non-empty region in the $(\beta,\lambda)$ plane, where only the diagonal solution $\bm{\sigma}^{(1,2,3)}$ is stable – but, of course, there could be other “spurious” states that can be stable in this region, making the disentanglement less efficient.
The saddle points of the free energy are stable when they are local minima of the RS free energy $f^{RS}=-\beta\mathcal{A}^{RS}$. This condition reads
$D_{\mu\nu}^{ab}=\frac{\partial^{2}f_{RS}}{\partial\overline{m}_{\mu}^{a}\partial\overline{m}_{\nu}^{b}}>0,$ (27)
that is, the Hessian matrix $D_{\mu\nu}^{ab}$ has to be positive definite.
These second-order derivative reads
$D_{\mu\nu}^{ab}=g_{ab}\delta_{\mu\nu}-\beta\sum_{c}g_{cb}g_{ca}\mathbb{E}_{\bm{\xi}}\left\{\xi^{\mu}\xi^{\nu}\left[1-\tanh^{2}\left(\beta\sum_{\rho\leq L}\xi^{\rho}\sum_{d}g_{cd}\overline{m}_{\rho}^{d}\right)\right]\right\}.$ (28)
In this expression we recognize the overlap $\overline{q}^{a}_{12}$ between the $(1,2)$ replicas in the same layer $a$, see (11), and the quantity $Q^{\mu\nu}_{a}$, defined as:
$\overline{q}^{a}_{12}=\mathbb{E}_{\bm{\xi}}\left\{\tanh^{2}\left(\beta\sum_{\rho\leq L}\xi^{\rho}\sum_{d}g_{ad}\overline{m}_{\rho}^{d}\right)\right\},$ (29)
$Q^{\mu\nu}_{a}=\mathbb{E}_{\bm{\xi}}\left\{\xi^{\mu}\xi^{\nu}\tanh^{2}\left(\beta\sum_{\rho\leq L}\xi^{\rho}\sum_{d}g_{ad}\overline{m}_{\rho}^{d}\right)\right\}.$ (30)
$D_{\mu\nu}^{aa}=\delta_{\mu\nu}\left[1-\beta\left(1-\overline{q}^{a}_{12}\right)+\lambda^{2}\sum_{c\neq a}(1-\overline{q}^{c}_{12})\right]+(1-\delta_{\mu\nu})\beta\left[Q^{\mu\nu}_{a}+\lambda^{2}\sum_{c\neq a}Q^{\mu\nu}_{c}\right],$
and the off-diagonal entries ($a\neq b$) as
$D_{\mu\nu}^{ab}=\delta_{\mu\nu}\lambda\left[-1+\beta(2-\overline{q}^{a}_{12}-\overline{q}^{b}_{12})+\lambda\sum_{c\neq a,b}(1-\overline{q}^{c}_{12})\right]+(1-\delta_{\mu\nu})\beta\lambda\left[-\left(Q^{\mu\nu}_{a}+Q^{\mu\nu}_{b}\right)+\lambda\sum_{c\neq a,b}Q^{\mu\nu}_{c}\right].$
Notice that, for $\mu=\nu$, $Q_{a}=\overline{q}^{a}_{12}$, while for $\mu\neq\nu$, $Q^{\mu\nu}_{a}$ is independent of the indices $\mu,\nu$ and it can be simply written as $Q_{a}=\mathbb{E}_{\bm{\xi}}\left\{\xi^{1}\xi^{2}\tanh^{2}\left(\beta\sum_{\rho\leq L}\xi^{\rho}\sum_{d}g_{ad}m_{\rho d}\right)\right\}$.
$t_{1}=1-\beta(1-\overline{q}^{a}_{12})-\beta\lambda^{2}\sum_{c\neq a}(1-\overline{q}^{c}_{12}),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \textrm{mult.}=K-L$ (31)
[…]
$t_{3}=t_{1}-(L-1)\beta Q_{a}-(L-1)\beta\lambda^{2}\sum_{c\neq a}Q_{c},\leavevmode\nobreak\ \,\,\textrm{mult.}=L-1.$ (33)
@@
The eigenvalues can be computed numerically for different values of $\beta,\lambda$ and for the related estimates of the magnetisation matrices $\bar{\bm{m}}^{(1,2,3)}$ and $\bar{\bm{m}}^{(h)}$. The stability of the saddle-point solutions depends on the sign of the smallest eigenvalue: if it is positive the solution is a minimum of the free energy $f_{RS}$ and therefore is said to be stable; otherwise, if negative, the solution is a saddle point or a maximum, and it is said to be unstable.
@@
The stability lines for the configurations $\bm{\sigma}^{(1,2,3)}$ and $\bm{\sigma}^{(h)}$ are reported in Fig. 4. It is worth stressing that there exists a non-vanishing region, where $\bm{\sigma}^{(h)}$ is unstable while $\bm{\sigma}^{(1,2,3)}$ is stable and the existence of such a region is a strictly necessary condition for this model to work. In fact, by initializing the system in $\bm{\sigma}^{(h)}$, we first want to move away from that state and eventually reach $\bm{\sigma}^{(1,2,3)}$. Consistently with the analysis led in Sec. 3.1, for this to occur the noise must be strictly positive. We also emphasize that the region determined here constitutes only an upper-bound as the instability and stability of, respectively, $\bm{\sigma}^{(h)}$ and $\bm{\sigma}^{(1,2,3)}$ do not directly imply that the former belongs to the attraction basin of the latter, that is, along its evolution, the system may bump into other stable states and remain nearby.
@@
Figure 4: Both panels present the range in the parameter space $(\beta,\lambda,H=0)$ where the three-layer model is expected to work as pattern disentangler. Below the red line the target configuration $\bm{\sigma}^{(1,2,3)}$ is stable, while above the green line the spurious configuration $\bm{\sigma}^{(h)}$ is unstable. The two lines are found by studying the sign of the Hessian $D_{\mu\nu}^{aa}$, obtained for $N\to\infty$ and $\gamma=0$, as reported in Sec. 3.2. The dashed lines are found by solving the self-consistency equations (11), by the fixed-point iteration method, starting from $\bm{\sigma}^{(h)}$, as explained in Sec. 3.3. More precisely, in the region between the two dashed curves, the solution found in this way corresponds to $\bm{\sigma}^{(1,2,3)}$, therefore in that region we expect that the machine can successfully work. Notice that the region determined by this method is, consistently, within the region outlined by stability analysis and, since it is derived from the self-consistency equations holding under the RS assumption and in the thermodynamics limit, it is expected to be subject to the same conditions. As a final test, useful to check possible finite-size corrections, we run MC simulations with a network made of $N=5000$ neurons and $K=5$ patterns, by initializing the system in the configuration $\bm{\sigma}^{(h)}$, updating it according to (13), and keeping track of whether the stable state corresponds or, still, it is strongly correlated with, $\bm{\sigma}^{(1,2,3)}$: if the experimental magnitudes $m_{1}^{1}$, $m_{2}^{2}$, and $m_{3}^{3}$ (or suitable permutations) are simultaneously larger than $0.99$ (left panel) or than 0.95 (right panel), the experiment is considered successful. Such trial is repeated $50$ times, for several choices of the parameters $\beta$ and $\lambda$, estimating the accuracy as the fraction of successful trails versus the number of trials (see the colormap).
@@
3.3 Numerical solutions of the saddle-point equations
@@
In this section we present results stemming from the numerical solution of the self-consistency equations (11). Before proceeding, a procedural remark is in order. In fact, for classical retrieval tasks, checking that the retrieval configuration is a solution of the saddle-point equation with a finite attraction basin, namely checking that it is a (local) minimum for the free-energy, is enough to state that the machine performs pattern retrieval. This can be inspected by solving the saddle-point equation via the fixed-point iteration method, starting from a configuration “close” to the retrieval one. On the other hand, this kind of procedure is not sufficient for the current task, that is, checking that the configuration $\bm{\sigma}^{(1,...,L)}$ is a (local) minimum for the free-energy is only a necessary condition here. Indeed, we need to require a stronger condition, namely, that the input configuration $\bm{\sigma}^{(h)}$ is unstable and belongs to the attraction basin of $\bm{\sigma}^{(1,...,L)}$. A possible way to check this is by looking for the solution of the saddle-point equation when the configuration $\bm{\sigma}^{(h)}$ is chosen as the starting point of the iterative method. Then, if that configuration constitutes a free-energy minimum, the fixed-point method will return $\bm{\sigma}^{*}=\bm{\sigma}^{(h)}$, otherwise, we expect that it will return the closest minimum, where the system is likely to end up.
@@
As mentioned in Sec. 2, the self-consistency equations (11) are rather awkward and their numerical solution, following the protocol described above, is computationally demanding. Thus, we will focus on the low-load regime, where, under the simplifying assumption $\gamma=0$, more friendly expressions can be recovered, as detailed in App. B. The numerical solution of these self-consistency equations, setting $L=3$, is plotted in Fig. 4 and in Fig. 5 for different choices of $\beta$, $\lambda$ and $H$, and compared with the results obtained by studying the stability of $\bm{\sigma}^{(h)}$ and of $\bm{\sigma}^{(1,2,3)}$ (see the previous Sec. 3.2) and with MC simulations (see the next Sec. 3.4). In particular, as $H$ gets larger, the successful region outlined by this method shrinks and moves toward larger values of $\lambda$ and smaller values of $\beta$, in fact, as $H$ gets larger the stability of the input configuration is reinforced, thus one needs a stronger inter-layer contribution and a higher degree of noise to destabilize it.
@@
Figure 5: We estimate the region in plane $(\beta,\lambda)$, where the three-layer model is expected to successfully disentangle mixtures of three patterns by solving the self-consistent equations (11) (dashed lines) and by running MC simulations (color map), in analogy to Fig. 4; in both cases we considered several values of the external field $H=0.0$ (left column), $H=0.1$ (middle column), $H=0.2$ (right column), and two different thresholds on the magnetizations $m>0.95$ (upper row), $m>0.99$ (lower row). For the first method, we set $\gamma=0$ and, as explained in Sec. 3.2, we found a region, bounded by the dashed lines, where the input configuration $\bm{\sigma}^{(h)}$ is attracted by the target output configuration $\bm{\sigma}^{(1,2,3)}$, thus within that region the system is expected to accomplish pattern disentanglement. For the second method, we set $N=5000$ and $K=5$, we initialize the system in the configuration $\bm{\sigma}^{(h)}$ and run the noisy dynamics (13) up to convergence to equilibrium. Then, the magnetizations of the three layers versus the patterns $\bm{\xi}^{1}$, $\bm{\xi}^{2}$, $\bm{\xi}^{3}$, are evaluated and if each of the three patterns is retrieved with a quality at least equal to the given threshold (no matter which layer retrieves a certain pattern), the simulation is considered as successful. The accuracy is finally evaluated over the sample of $50$ trials and represented by the color map.
3.4 Monte Carlo simulations
@@
After the previous theoretically-driven analysis, we now tackle the problem computationally as this allows us to corroborate the former which is subject to the RS and the thermodynamic-limit assumptions. Moreover, the previous theoretically-driven analysis only provided an upper-bound for the region in the space $(\beta,\lambda,H)$ where we can expect the machine to work, without quantifying how well and how likely the machine can work. To answer this question, here, we initialize the system in the spurious state $\bm{\sigma}^{(h)}$, we let it evolve according to (13) and, once a stable state is reached, we check whether this is retrieving the single components, that is, if it corresponds to $\bm{\sigma}^{(1,2,3)}$ (or any suitable permutation). We repeat the experiment several times, counting the number of successful experiments, where “successful” means that the magnitudes of the observed magnetizations $m_{1}^{1},m_{2}^{2},m_{3}^{3}$ are larger than a certain threshold. Finally, the accuracy is evaluated as the fraction between the number of successful experiments and the overall number of experiments, and plotted in Fig. 4 and in Fig. 5. Remarkably, there exists a region, inside the upper-bound determined analytically, where the accuracy is unitary or very close to one, and the existence of such a region guarantees that the machine can disentangle the inputted spurious state. Of course, this region gets wider as the threshold for success is lowered.
In the end, the time evolution of the magnetizations $m_{1}^{1},m_{2}^{2}$, and $m_{3}^{3}$ is shown in Fig. 6, for different choices of the parameters, where, according to what shown in Fig. 5 we expect an ergodic phase (upper panels), a correct disentanglement (middle panels), and a deadlock in the input state (lower panels). The consistency between theoretical and computational results is fully recovered.
@@
Figure 6: These plots show the evolution of the Mattis magnetizations $m_{\mu}^{a}$ for $\mu=1,...,K$ (different labels correspond to different colors) and for $a=1,2,3$ (different layers correspond to different columns) versus the number of MC steps – one MC step corresponds to $N$ random extractions of the index $i\in\{1,...,N\}$ that identifies the neuron to be updated according to the rule (13). More precisely, here we set $N=5000$, $K=50$ $H=0.2$ and $\lambda=0.2$, while different values of $\beta$ are chosen: $\beta=1$ (upper row), $\beta=2$ (middle row), $\beta=3$ (lower row); in agreement with the findings presented in Fig. 4, the emerging behavior is, respectively, ergodic, disentangled, and stuck in the spurious state.
Figure 7: We evaluate $\mathcal{H}(\bm{\sigma};\lambda,H,\bm{\xi},\bm{h})/N$ (left panel) and ~$\widetilde{\mathcal{H}}(\bm{\sigma};\lambda,H,\bm{\xi},\bm{h})/N$ at the configurations $\bm{\sigma}^{(1,2,3)}$, $\bm{\sigma}^{(1,1,1)}$, $\bm{\sigma}^{(1,1,1^{\prime})}$, and $\bm{\sigma}^{(h)}$ according to eqs. (16),(20), (23) and (35)-(38), and we keep track of the configurations displaying the lower energy as the parameters $H$ and $\lambda$ are varied. For the first model the region where the target configuration is energetically favoured is the one corresponding to relatively large values of $\lambda$ and relatively small values of $H$, while for the second model that region encompasses the whole region below the curve $H=1/2+2\lambda(3/4+\gamma)^{2}$. Different values of $\gamma$ are also considered: $\gamma=0.1$ (solid line), $\gamma=0.05$ (dashed line), and $\gamma=0.005$ (dotted line): the arrows point in the direction of increasing $\gamma$.
Figure 8: We consider the system described by the revised Hamiltonian (34) and we simulate its evolution starting from the configuration $\bm{\sigma}^{(h)}$ and iteratively applying the noisy dynamics (13), up to convergence to equilibrium. Analogously with what done in Figs. 4 - 5, we set $N=5000$ and $K=50$ and we repeated the MC simulation $50$ times for each sampled point of the ($\lambda$, $\beta^{-1}$) plane and for various values of the external field $H=0.0$ (first row), $H=0.2$ (second row), $H=0.4$ (third row), $H=0.6$ (fourth row); next, the magnetizations of the three layers versus the patterns $\bm{\xi}^{1}$, $\bm{\xi}^{2}$, $\bm{\xi}^{3}$, are evaluated and, if each of the three patterns is retrieved with a quality at least equal to the given threshold, the simulation is considered as successful. The accuracy, represented by the color map, is then evaluated over the sample of $50$ trials. Finally, notice that, unlike Figs. 4-5 here we plotted data versus $\beta^{-1}$ to highlight that the system is able to accomplish the task even in the noiseless case $\beta^{-1}\to 0$.
4 A performance-driven revision
The analysis carried on in the previous sections showed that an assembly of interacting Hopfield networks is able to accomplish tasks that are not achievable by a single Hopfield network. However, since the preliminary results presented in Sec. 3.1, one could realize that this model is probably not the optimal one if specifically interested in pattern disentanglement; indeed, our purpose is the investigation of non-trivial phenomena emerging from the interaction of networks, rather than specifically pattern disentanglement, see Jean2024 . In fact, our target configuration is not a ground state for the model and, as $\beta\to\infty$, the system would remain stuck in the input configuration. We recognize that the intra-layer interactions work properly by favoring the alignment of each layer to patterns, on the other hand, the inter-layer interactions, which should inhibit the retrieval of the same pattern by different layers, tend to favor the staggered configuration instead of the target configuration. This flaw can be fixed by revising the coupling between different layers. Indeed, this term explicitly breaks the layer-wise spin-flip symmetry of our model and stabilizes the state $\bm{\sigma}^{(1,1,1^{\prime})}=(\bm{\xi}^{1},\bm{\xi}^{1},-\bm{\xi}^{1})$, which is among the states that most significantly hinder the network’s disentanglement task. A modified Hamiltonian reads as:
~$\begin{array}[]{lll}\widetilde{\mathcal{H}}(\bm{\sigma};\lambda,H,\bm{\xi},\bm{h})=-N\displaystyle\sum\limits_{\mu=1}^{K}\sum_{a=1}^{L}(m^{a}_{\mu})^{2}-H\displaystyle\sum\limits_{i=1}^{N}\sum_{a=1}^{L}h^{a}_{i}\sigma_{i}^{a}+N\lambda\sum_{\begin{subarray}{c}a,b=1\\ a\neq b\end{subarray}}^{L}\left(\displaystyle\sum\limits_{\mu=1}^{K}m^{a}_{\mu}m^{b}_{\mu}\right)^{2}\end{array}$ (34)
and it differs from the original one (5) only in the last contribution in the right-hand side of (34), which now features a quadratic sum over the heterogeneous product of magnetizations, rather than a linear one. This modification has two advantages: first, in the absence of an external field (i.e., $H=0$), it makes the Hamiltonian invariant under layer-wise spin-flip, further, it inhibits the relaxation towards states like $\bm{\sigma}^{(1,1,1’)}$, making, as we will see, the disentanglement task more robust and stable even at very low noise.
~$\frac{\widetilde{\mathcal{H}}(\bm{\sigma}^{(1,2,3)})}{N}\underset{c.l.t.}{\sim}-3(1+\gamma)-\frac{3}{2}H+x\frac{\tilde{\mathcal{C}}^{(1,2,3)}}{\sqrt{N}},$ (35)
~$\frac{\widetilde{\mathcal{H}}(\bm{\sigma}^{(1,1,1)})}{N}\underset{c.l.t.}{\sim}-3(1+\gamma)+3\lambda(1+\gamma)^{2}-\dfrac{3}{2}H+x\frac{\tilde{\mathcal{C}}^{(1,1,1)}}{\sqrt{N}},$ (36)
~$\frac{\widetilde{\mathcal{H}}(\bm{\sigma}^{(1,1,1^{\prime})})}{N}\underset{c.l.t.}{\sim}-3(1+\gamma)+3\lambda(1+\gamma)^{2}-\dfrac{1}{2}H+x\frac{\tilde{\mathcal{C}}^{(1,1,1^{\prime})}}{\sqrt{N}}$ (37)
~$\frac{\widetilde{\mathcal{H}}(\bm{\sigma}^{(h)})}{N}\underset{c.l.t.}{\sim}-3\left(\dfrac{3}{4}+\gamma\right)+3\lambda\left(\dfrac{3}{4}+\gamma\right)^{2}-3H+x\frac{\tilde{\mathcal{C}}^{(h)}}{\sqrt{N}}.$ (38)
By comparison with eqs. (16), (18), (20), and (23), we see that (~$\widetilde{\mathcal{H}(}\bm{\sigma}^{(1,2,3)})$ is asymptotically the same as ${\mathcal{H}(\bm{\sigma}^{(1,2,3)})}$, moreover, now $\lambda$ has a stronger effect in making the configuration $\bm{\sigma}^{(1,1,1)}$ unstable and its influence on $\bm{\sigma}^{(1,1,1^{\prime})}$ shifts from positive to negative; as for $\bm{\sigma}^{(h)}$, this state is slightly favored in the current setting, especially for low loads. As a result, here, for $H$ relatively small, $\bm{\sigma}^{(1,2,3)}$ is always prevailing over $\bm{\sigma}^{(1,1,1^{\prime})}$, see Fig. 7.
Finally, MC simulations analogous to those presented in Sec. 3.4 have been run for the system described by the Hamiltonian (34) and for various parameter settings. The results, presented in Fig. 8, show that the region where the spurious-state disentanglement occurs successfully is no longer vanishing in the zero-temperature limit. Furthermore, when the temperature increases (e.g., at $\beta=2$), the region of high accuracy performance is significantly enlarged compared to the results obtained with the Hamiltonian (3) and presented in Fig. 5. The robustness of these results is checked in Fig. 9, where we executed a numerical test with a nonrandom data set, where the patterns represent digits and their mixture (see the left-most panels in the figure) is used as input for a three-layer network where neurons interact according to (34).

### v2

@@ 4.2 Stability analysis in the low-load, noisy, and zero-field regime
Figure 3: The solid lines represent the numerical solution of the self-consistency equations (11) in the low-load regime and in the absence of external field, obtained by applying the fixed-point iteration method with initial point given by $\bar{\bm{m}}_{\{\mu\leq L\}}^{(1,2,3)}$ (left) and by $\bar{\bm{m}}_{\{\mu\leq L\}}^{(h)}$ (right), see eqs. (16)-(17). These numerical solutions preserve the structure of the initial datum, specifically, on the left, the solid lines show the behavior of $\bar{m}_{1}^{1}=\bar{m}_{2}^{2}=\bar{m}_{3}^{3}$ while $\bar{m}_{\mu}^{a}$ is vanishing for $\mu\neq a$; on the right, the the solid lines show the behavior of $\bar{m}_{\mu}^{a}$, that coincides for any $a\in[1,2,3]$ and $\mu\in[1,2,3]$. The persistency in the structure of the solution is lost at a certain value of $\beta^{-1}$, highlighted by the vertical dotted lines: beyond these values, that depend on $\lambda$ (see the common legend on the right), solutions with a different structure appear, and these correspond, for instance, to the state $\bm{\sigma}^{(1,1,1^{\prime})}$.
As already recalled, the solutions of the self-consistency equations (11) are not necessarily equilibrium states as we also need to check that these extremal states are minima of the free energy $f=-\beta\mathcal{A}$. We now proceed in this direction and we denote with
$D_{\mu\nu}^{ab}=\frac{\partial^{2}f_{RS}}{\partial\overline{m}_{\mu}^{a}\partial\overline{m}_{\nu}^{b}}$ (18)
the entries of the Hessian matrix related to the free energy (8), recalling that here $\gamma=0$. Next, we determine the conditions on the control parameters $\beta$ and $\lambda$ under which the Hessian matrix is positive definite when evaluated at the magnetizations $\bar{\bm{m}}^{(1,2,3)}$ and $\bar{\bm{m}}^{(h)}$.
Starting from the second-order derivative
$D_{\mu\nu}^{ab}=g_{ab}\delta_{\mu\nu}-\beta\sum_{c=1}^{L}g_{cb}g_{ca}\mathbb{E}_{\bm{\xi}}\left\{\xi^{\mu}\xi^{\nu}\left[1-\tanh^{2}\left(\beta\sum_{\rho=1}^{L}\xi^{\rho}\sum_{d=1}^{L}g_{cd}\overline{m}_{\rho}^{d}\right)\right]\right\}$ (19)
@@
and, following some straightforward algebraic manipulations (see App. D) we get the spectrum of the Hessian matrix. The stability of an extremal state depends on the sign of the smallest eigenvalue: if it is positive the solution is a minimum of the free energy $f_{RS}$ and therefore is said to be stable; otherwise, if negative, the solution is a saddle point or a maximum, and it is said to be unstable.
@@
The stability lines for the configurations $\bm{\sigma}^{(1,2,3)}$ and $\bm{\sigma}^{(h)}$ are reported in Figure 4. It is worth stressing that there exists a non-vanishing region, where $\bm{\sigma}^{(h)}$ is unstable while $\bm{\sigma}^{(1,2,3)}$ is stable and the existence of such a region is a strictly necessary condition for this model to work. In fact, by initializing the system in $\bm{\sigma}^{(h)}$, we first want to move away from that state and eventually reach $\bm{\sigma}^{(1,2,3)}$ – but, of course, there could be other “spurious” states that can be stable in this region, making the disentanglement less efficient. Consistently with the analysis led in Sec. 4.1, for this to occur the noise must be strictly positive. We also emphasize that the region determined here constitutes only an upper-bound as the instability and stability of, respectively, $\bm{\sigma}^{(h)}$ and $\bm{\sigma}^{(1,2,3)}$ do not directly imply that the former belongs to the attraction basin of the latter, that is, along its evolution, the system may bump into other stable states and remain nearby.
@@
Figure 4: Both panels present the range in the parameter space $(\beta,\lambda,H=0)$ where the three-layer model is expected to work as pattern disentangler. Below the red line the target configuration $\bm{\sigma}^{(1,2,3)}$ is stable, while above the green line the spurious configuration $\bm{\sigma}^{(h)}$ is unstable. The two lines are found by studying the sign of the Hessian $D_{\mu\nu}^{aa}$, obtained for $N\to\infty$ and $\gamma=0$, as reported in Sec. 4.2 and App. D. The dashed lines are found by solving the self-consistency equations (11), by the fixed-point iteration method, starting from $\bm{\sigma}^{(h)}$, as explained in Sec. 4.3. More precisely, in the region between the two dashed curves, the solution found in this way corresponds to $\bm{\sigma}^{(1,2,3)}$, therefore in that region we expect that the machine can successfully work. Notice that the region determined by this method is, consistently, within the region outlined by stability analysis and, since it is derived from the self-consistency equations holding under the RS assumption and in the thermodynamics limit, it is expected to be subject to the same conditions. As a final test, useful to check possible finite-size corrections, we run MC simulations with a network made of $N=5000$ neurons and $K=5$ patterns, by initializing the system in the configuration $\bm{\sigma}^{(h)}$, updating it according to (13), and keeping track of whether the stable state corresponds or, still, it is strongly correlated with, $\bm{\sigma}^{(1,2,3)}$: if the experimental magnitudes $m_{1}^{1}$, $m_{2}^{2}$, and $m_{3}^{3}$ (or suitable permutations) are simultaneously larger than $0.99$ (left panel) or than 0.95 (right panel), the experiment is considered successful. Such trial is repeated $50$ times, for several choices of the parameters $\beta$ and $\lambda$, estimating the accuracy as the fraction of successful trails versus the number of trials (see the colormap). We remark that an overall very good agreement among the theoretical predictions and the numerical outcomes is obtained.
@@
4.3 Checking disentanglement properties by numerical solutions of the saddle-point equations
@@
For classical retrieval tasks, checking that the retrieval configuration is a solution of the saddle-point equation with a finite attraction basin, namely checking that it is a (local) minimum for the free-energy, is enough to state that the machine performs pattern retrieval. This can be inspected by solving the saddle-point equation via the fixed-point iteration method, starting from a configuration “close” to the retrieval one, as previously done in Sec. 4.2. On the other hand, this kind of procedure is not sufficient for the current task, that is, checking that the configuration $\bm{\sigma}^{(1,...,L)}$ is a (local) minimum for the free-energy is only a necessary condition here. Indeed, we need to require a stronger condition, namely, that the input configuration $\bm{\sigma}^{(h)}$ is unstable and belongs to the attraction basin of $\bm{\sigma}^{(1,...,L)}$. A possible way to check this is by looking for the solution of the saddle-point equation when the configuration $\bm{\sigma}^{(h)}$ is chosen as the starting point of the iterative method. Then, if that configuration constitutes a free-energy minimum, the fixed-point method will return $\bm{\sigma}^{*}=\bm{\sigma}^{(h)}$, otherwise, we expect that it will return the closest minimum, where the system is likely to end up.
@@
As mentioned in Sec. 2, the self-consistency equations (11) are rather awkward and their numerical solution, following the protocol described above, is computationally demanding. Thus, we will focus on the low-load regime, where, under the simplifying assumption $\gamma=0$, more friendly expressions can be recovered, as detailed in App. B. The numerical solution of these self-consistency equations, setting $L=3$, is plotted in Figure 4 and in Figure 5 for different choices of $\beta$, $\lambda$ and $H$, and compared with the results obtained by studying the stability of $\bm{\sigma}^{(h)}$ and of $\bm{\sigma}^{(1,2,3)}$ (see the previous Sec. 4.2) and with MC simulations (see the next Sec. 4.4). In particular, as $H$ gets larger, the successful region outlined by this method shrinks and moves toward larger values of $\lambda$ and smaller values of $\beta$, in fact, as $H$ gets larger the stability of the input configuration is reinforced, thus one needs a stronger inter-layer contribution and a higher degree of noise to destabilize it.
@@
Figure 5: We estimate the region in plane $(\beta,\lambda)$, where the three-layer model is expected to successfully disentangle mixtures of three patterns by solving the self-consistent equations (11) (dashed lines) and by running MC simulations (color map), in analogy to Figure 4; in both cases we considered several values of the external field $H=0.0$ (left column), $H=0.1$ (middle column), $H=0.2$ (right column), and two different thresholds on the magnetizations $m>0.95$ (upper row), $m>0.99$ (lower row). For the first method, we set $\gamma=0$ and, as explained in Sec. 4.3, we found a region, bounded by the dashed lines, where the input configuration $\bm{\sigma}^{(h)}$ is attracted by the target output configuration $\bm{\sigma}^{(1,2,3)}$, thus within that region the system is expected to accomplish pattern disentanglement. For the second method, we set $N=5000$ and $K=5$, we initialize the system in the configuration $\bm{\sigma}^{(h)}$ and run the noisy dynamics (13) up to convergence to a stationary state. Then, the magnetizations of the three layers versus the patterns $\bm{\xi}^{1}$, $\bm{\xi}^{2}$, $\bm{\xi}^{3}$, are evaluated and if each of the three patterns is retrieved with a quality at least equal to the given threshold (no matter which layer retrieves a certain pattern), the disentanglement achieved in that simulation is considered as successful. The accuracy is finally evaluated over the sample of $50$ trials and represented by the color map.
4.4 Checking disentanglement properties by Monte Carlo simulations
After the previous theoretically-driven analysis, we now tackle the problem computationally as this allows us to corroborate the theory, which is subjected to the RS and the thermodynamic limit assumptions. Moreover, the previous theoretically-driven analysis only provided an upper-bound for the region in the space $(\beta,\lambda,H)$ where we can expect the machine to work, without quantifying how well and how likely the machine can work. Here, to answer this question, we run MC simulations, whose details, along with pseudo codes and a time consumption analysis are presented in App. E.
@@ 4.4 Checking disentanglement properties by Monte Carlo simulations
In our experiments we initialize the system in the spurious state $\bm{\sigma}^{(h)}$, we let it evolve according to (13) and, once a stable state is reached, we check whether this is retrieving the single components, that is, if it corresponds to $\bm{\sigma}^{(1,2,3)}$ (or any suitable permutation): , see Figure 6 where we inspect the time evolution of the magnetizations $m_{1}^{1},m_{2}^{2}$, and $m_{3}^{3}$. We repeat the experiment several times, spanning over the parameters $\beta,\lambda,H$ and counting the number of successful experiments, where “successful” means that the magnitudes of the observed magnetizations $m_{1}^{1},m_{2}^{2},m_{3}^{3}$ are larger than a certain threshold. Finally, the accuracy is evaluated as the fraction between the number of successful experiments and the overall number of experiments, and plotted in Figure 4 and in Figure 5. Remarkably, there exists a region, inside the upper-bound determined analytically, where the accuracy is unitary or very close to one, and the existence of such a region guarantees that the machine can disentangle the inputted spurious state. Of course, this region gets wider as the threshold for success is lowered.
The robustness of these results and their scalability versus $L$ is discussed in App. F, where the analysis for the case $L=5$ are reported.
@@
Figure 6: These plots show the evolution of the Mattis magnetizations $m_{\mu}^{a}$ for $\mu=1,...,K$ (different labels correspond to different colors) and for $a=1,2,3$ (different layers correspond to different columns) versus the number of MC steps – one MC step corresponds to NN random extractions of the index $i\in\{1,...,N\}$ that identifies the neuron to be updated according to the rule (13), see also App. E. More precisely, here we set $N=5000$, $K=50$ $H=0.2$ and $\lambda=0.2$, while different values of $\beta$ are chosen: $\beta=1$ (upper row), $\beta=2$ (middle row), $\beta=3$ (lower row); in agreement with the findings presented in Figure 4, the emerging behavior is, respectively, ergodic, disentangled, and stuck in the spurious state.
We close this section by noting that, as we move away from the Rademacher dataset toward more structured patterns, the network’s performance is expected to deteriorate. This is because Hebbian couplings are known to be particularly effective when the stored memories are (approximately) orthogonal. It is therefore worth considering a benchmark dataset to verify whether the disentanglement capabilities of the model (5) are preserved. In Figure 7, we provide numerical evidence supporting this. However, we emphasize that in this case the performance is more sensitive to parameter tuning, and the successful region in the $(\beta,\lambda)$ plane is expected to be smaller than in the Rademacher case [^11] Also, in general, the optimal parameter setting might depend on the mixed patterns.. This observation further underscores the importance of having a solid theoretical foundation. We also note, consistently with the analytical results presented earlier (see, e.g., Sec. 4.1), that a certain degree of noise (i.e., $\beta^{-1}>0$) is still required, implying that the retrieval of the deconvolved patterns exhibits some imperfections. To enable effective performance even in the noiseless regime, certain adjustments can be made — specifically, by introducing higher-order inter-layer interactions, as discussed in App.G.